# Controllable Generation with Text-to-Image Diffusion Models: A Survey

### v1

@@ 4.2.1 Spatial-Conditional Score Prediction
While the above methods hope to generate images fully aligned to given conditions, some methods study employing coarse and incomplete spatial condition [115, 145, 146]. Specically, LooseControl [145] extracts proxy depth for 3D box control from images and finetunes the ControlNet [190] by LoRA [192], enabling to create complex environments (e.g., rooms, street views) by specifying only scene boundaries and locations of primary objects.
• Training-free Score Prediction. Since attention mechanism explicitly model the relationships between text and image tokens, modulating attention map becomes a pivotal training-free technique for controlling structure in score prediction [26, 137, 135, 142, 118]. eDiff-I [26] presents a techinique named “paint-with-words” (also known as pww), rectifying the cross-attention maps of each word by the correspondence segmentation maps to control the location of objects. Additionally, DenseDiffusion [137] introduces a more extensive modulation method by devising multiple regularization, enhancing the precision and flexibility of layout control in score prediction.
4.2.2 Spatial-Guided Score Estimation
While numerous methods adhere to the paradigm of conditional score prediction like ControlNet, some studies have explored spatial controlling through sptail-guided score estimation [121, 124, 131, 127]. Notably, LGP [121] stands as an early pioneer, which innovatively introduces a Latent Edge Predictor, designed to extrapolate sketch information from a series of intermediate features within a UNet architecture. It employs the degree of similarity between condition sketch and predicted sketch to compute gradients, which are then utilized to guide the score estimation process. Its methodologies and insights have been a source of inspiration for numerous subsequent research endeavors in this field[131, 133, 134, 220]. ZestGuide [134] leverages segmentation maps extracted from cross-attention layers, aligning generation with input masks through gradient-based guidance during denoising.
4.3 Advanced Text-Conditioned Generation
While text serves as the fundamental condition in text-to-image diffusion models, several challenges persist in this domain. First, text-guided synthesis, particularly with complex text involving multiple subjects or enriched descriptions, often encounters issues of textual misalignment. Furthermore, the predominant training of these models on English datasets has led to a notable lack in multilingual generation capabilities. To address this limitation, innovative approaches aimed at expanding the linguistic scope of these models have been proposed.
• Improving Textual Alignment. Textual alignment plays a pivotal role in text-to-image diffusion models, providing essential control over the generation process. Despite being trained on multimodal text-image datasets, these generative models often struggle to precisely capture and reflect the full spectrum of information contained in textual descriptions. To address this challenge, various innovative approaches have been developed [150, 149, 152]. Specifically, Attend-and-Excite [150] represents an early effort in this area, introducing an attention-based Generative Semantic Nursing (GSN) mechanism. This mechanism refines cross-attention units to more effectively ensure that all subjects described in the text prompt are accurately generated. Structure Diffusion [149] employs linguistic insights to manipulate the cross-attention map, aiming for more accurate attribute binding and improved image composition. Ge et al. [152] propose a rich-text-to-image framework, which initially processes plain text through a diffusion model to gather attention maps, noised generation, and residual feature maps. Subsequently, rich texts are formatted as JSON to provide detailed attributes for each token span, enhancing the model’s capacity to align visual content with complex textual descriptions. Additionally, SynGen [153] employs a unique methodology in text-to-image generation by first conducting a syntactic analysis of the text prompt. This analysis aims to identify entities and their modifiers within the prompt. Following this, SynGen utilizes a novel loss function designed to align the cross-attention maps with the linguistic bindings as indicated by the syntax. Furthermore, Tailored Visions [154] leverages historical user interactions with the system to rewrite user prompts to enhance the expressiveness and alignment of user prompts with their intended visual outputs. To improve textual alignment of a long paragraph (up to 512 words), Wu et al. [155] introduce an informative-enriched diffusion model for paragraph-to-image generation task, termed ParaDiffusion, which employ a large language model (e.g., Llama V2 [221]) to encode long-form text, followed by fine-tuning with LoRA [192] to align text-iamge feature spaces in generation.
@@ 4.3 Advanced Text-Conditioned Generation
• Multilingual-Guided Generation. GlueGen [151] aligns multilingual language model (e.g., XLM-Roberta [222]) with existing text-to-image models, allowing for the generation of high-quality images from captions beyond English. PEA-Diffusion [156] is a proposed simple plug-and-play language transfer method based on knowledge distillation, where a lightweight MLP-like parameter efficient adapter with only 6M parameters is trained under teacher knowledge distillation along with a small parallel data corpus.
4.4 In-Context Generation
The in-context generation task involves understanding and performing specific tasks on new query images based on a pair of task-specific example images and text guidance.
@@ 4.4 In-Context Generation
Wang et al. [157] introduced Prompt Diffusion, a novel approach that is jointly trained over multiple tasks using in-context prompts. This method has shown impressive results in high-quality in-context generation for trained tasks and effectively generalizes to new, unseen vision tasks with relevant prompts. Building upon this, Chen et al. [158] further enhance Prompt Diffusion by incorporating a vision encoder-modulated text encoder. This innovation addresses several challenges, including costly pre-training, restrictive problem formulations, limited visual comprehension, and insufficient generalizability to out-of-distribution tasks. Moreover, Najdenkoskapropose a novel framework that separates the encoding of the visual context and preserving the structure of the query images. This results in the ability to learn from the visual context and text prompts, but also from either one of them.
4.5 Brain-Guided Generation
@@
The brain-guided generation tasks focus on controlling image creation directly from brain activities, such as electroencephalogram (EEG) recordings and functional magnetic resonance imaging (fMRI), bypassing the need to translate thoughts into text. Early studies in this domain have employed Generative Adversarial Networks (GANs) and Variational Autoencoders (VAEs) to reconstruct visual images from brain signals [223, 224, 225, 226]. More recently, advancements have been made with the adoption of visual diffusion models, offering enhanced capabilities in accurately translating complex brain activities into coherent visual representations[159, 160, 161, 162, 163, 164, 165].
@@ 4.6 Sound-Guided Generation
GlueGen [151] aligns multi-modal encoders such as AudioCLIP with the Stable Diffusion model, enabling sound-to-image generation. Yang et al. [166] propose a unified framework “Align, Adapt, and Inject” (AAI) for sound-guided image generation, editing, and stylization. In particular, this method adapts input sound into a sound token, like an ordinary word, which can plug and play with existing powerful diffusion-based Text-to-Image models.
4.7 Text Rendering
The task of text rendering within synthesized images is pivotal, especially given the widespread application of text in various visual forms like posters, book covers, and memes.
Drawing inspiration from the analysis in unCLIP [38], which highlights the inadequacy of raw CLIP text embeddings in accurately modeling the spelling information in prompts, subsequent efforts such as eDiff-I [26] and Imagen [24] have sought to harness the capabilities of large language models like T5 [37], trained on text-only corpora, as text encoders in image generation. Additionally, DeepFloyd IF  [^2] https://github.com/deep-floyd/IF, following the design principles of Imagen [24], has demonstrated impressive proficiency in rendering legible text on images, showcasing a significant advancement in this challenging domain.
@@ 4.7 Text Rendering
Meanwhile, some approaches are designed to improve text rendering capability for existing text-to-image diffusion models [167, 168, 169, 171, 172, 173, 174]. Liu et al. [167] find popular text-to-image models lack character-level input features, making it much harder to predict a word’s visual makeup as a series of glyphs. GlyphControl [170] leverages additional glyph conditional information to enhance the performance of the off-the-shelf Stable-Diffusion model in generating accurate visual text. TextDiffuser et al. [169] first generates the layout of keywords extracted from text prompts and then generates images conditioned on the text prompt and the generated layout. The authors also contribute a large-scale text images dataset with OCR annotations, MARIO-10M, containing 10 million image-text pairs with text recognition, detection, and character-level segmentation annotations. Zhang et al. [174] proposed Diff-Text, a training-free scene text generation framework for any language. Diff-text leverages rendered sketch images as priors to render text by ControlNet [190] and propose a localized attention constraint to address the unreasonable position problem of scene text.
@@ 5 Controllable Generation with Multiple Conditions
The multi-condition generation task aims to generate images under multiple conditions, such as generating a specific person in a user-defined pose or generating people in three personalized identities. In this section, we conduct a comprehensive overview of these methods from a technical perspective, categorizing them into joint training (Section5.1), weight fusion (Section5.3), attention-based integration (Section5.4), guidance fusion (Section5.5), and continual learning (Section5.2). Note that some of the other controllable generation methods also demonstrate multi-condition synthesis capability without dedicated designs [41, 42, 215].
@@ 5.1 Joint Training
Designing a multi-condition framework and jointly training them is a simple yet effective route to realize multi-condition generation. These methods generally focus on multi-condition encoders and training strategies.
Composer [175] projects all conditions (including text caption, depthmap, sketch, and etc.) into uniform-dimensional embeddings with the same spatial size as the noisy latent using stacked convolutional layers. It leverages a joint training strategy to generate images from a set of representations, where it uses an independent dropout probability of 0.5 for each condition, a probability of 0.1 for dropping all conditions, and a probability of 0.1 for retaining all conditions. Additionally, Cocktail [176] proposes the controllable normalization method (ControlNorm), which has an additional layer to generate two sets of learnable parameters conditioned on all modalities. These two sets of parameters are used to fuse the external conditional signals and the original signals.
@@ 5.2 Continual Learning
[…]

### v2

@@ 3.1.2 Adapter-based Conditional Score Prediction
While many aim for strict adherence to provided controls, other methods support coarse or incomplete spatial inputs [336, 18, 160]. Specifically, LooseControl [18] extracts proxy depth from images to define 3D box controls and fine-tunes ControlNet [323] via LoRA [98], enabling the creation of complex environments (e.g., rooms, street scenes) by specifying only scene boundaries and key object locations.
• Image Personalization. Adapter-based image personalization methods employ encoders to embed concepts (e.g., subject, human identity, style), offering a significant speed advantage over tuning-based approaches when extracting concepts from images.
Some methods adopt a domain-agnostic strategy, training encoders on open-world images to extract generalized subject-level conditions [283, 174, 39, 4, 173, 113, 188, 241, 102, 54, 243, 201]. These methods typically leverage large pretrained encoders such as CLIP [208] and BLIP-2 [137], fine-tuning only lightweight projection layers [283, 174, 4]. ELITE [283], for example, integrates a global mapping network and a local mapping network based on CLIP [208]. The global network transforms hierarchical image features into multiple text embeddings, while the local network infuses patch features into cross-attention layers for detailed reconstruction. BLIP-Diffusion [136] advances customization by pre-training a BLIP-2 [137] encoder for text-aligned image representation and developing a task for learning subject representations, enabling the generation of novel subject renditions. Following on E4T [69], Arar et al. [4] introduce an encoder for acquiring text embeddings and propose a hypernetwork to predict LoRA-style attention weight offsets in UNet. SuTI [39] takes a unique approach inspired by apprenticeship learning [1], training a vast array of expert models on millions of internet image clusters. The apprentice model is then taught to imitate these experts’ behaviors. CAFE [344] build a customization assistant based on pre-trained large language model and diffusion model.
Some works design domain-aware encoders tailored to targeted domains [69, 232, 47]. In person-driven settings, facial images are encoded to provide identity conditions [307, 290, 257, 43, 37, 145, 222, 195, 275, 79, 268, 52, 233, 146, 70, 155]. Face0 [257] and DreamIdentity [43] employ pretrained face recognition encoders [23]; Face0 uses Inception-ResNet-V1 [248], while DreamIdentity introduces a ViT-style M2ID encoder [60]. Beyond multimodal and face-recognition encoders [23], $\mathscr{W}^{+}$ Adapter [140] and PreciseControl [191] leverage GAN inversion [20, 21] encoders as an alternative identity pathway.
Furthermore, some person-driven methods study the mechanism of combining textual embeddings with identity embeddings [290, 145, 52]. To balance identity preservation and editability, FastComposer [290] and PhotoMaker [145] fuse text prompts with visual features from reference images; specifically, FastComposer mixes human-related text tokens (e.g.,“man”, “woman”) with visual features via a multilayer perceptron, and PhotoMaker applies two MLP layers to fuse image embeddings with the corresponding human-related embedding, then updates the latter with the fused representation. Beyond an identity encoder, some works additionally employ a spatial condition encoder to improve generation quality [84]. T his task also benefits from face segmentation masks and skeletal cues obtained from off-the-shelf models or annotations [290, 145, 2, 194, 117]. For example, Stellar [2] uses face masks to remove background during preprocessing, sharpening focus on identity, while other methods leverage face masks to construct [290, 145, 194] or adjust [108] loss functions.
Notably, task-specific encoders can be effective for image personalization. For style-driven generation, several studies use VGG features to better capture low-level style [30, 202]. Prompt-free Diffusion [298] introduces SeeCoder, composed of a backbone encoder, a decoder, and a query transformer, enabling reference images to serve as conditions in lieu of text prompts.
• View-conditioned Generation. Adapter-based view-conditioned generation aims to leverage explicit geometric or panoramic constraints within generative models to ensure spatial consistency across different viewpoints [230, 11, 312, 317, 264, 129, 17, 190]. PreciseCam [17] converts four simple extrinsic and intrinsic camera parameters to a PF-US map and then controls generation by a trained ControlNet. StitchDiffusion [264] extends adapter-based controllable generation to the domain of 360-degree panoramas by fine-tuning a T2I diffusion model with LoRA and introducing a stitching-aware denoising strategy. This approach ensures seamless global geometry and strong generalization for panoramic scene synthesis.
• Advaned Text-conditioned Generation. To better extract faithful textual semantics, some works leverage LLM to replace the CLIP text encoder [45, 285, 251]. To improve the textual alignment of a long paragraph (up to 512 words), Wu et al. [285] introduce an informative-enriched diffusion model for paragraph-to-image generation task, termed ParaDiffusion, which employ a large language model (e.g., Llama V2 [254]) to encode long-form text, followed by fine-tuning with LoRA [98] to align text-iamge feature spaces in generation.
Other methods design an extra pipeline or textual encoders to improve textual controllability [45, 301]. Tailored Visions [45] introduces a prompt rewriting system, leveraging historical user interactions to rewrite user prompts to enhance the expressiveness and alignment of user prompts with their intended visual outputs.
@@
The text encoder is also studied to extend to multilingual version. GlueGen [205] aligns multilingual language models (e.g., XLM-Roberta [49]) with existing text-to-image models, allowing for the generation of high-quality images from captions beyond English. PEA-Diffusion [172] is a proposed simple plug-and-play language transfer method based on knowledge distillation, where a lightweight MLP-like parameter-wefficient adapter with only 6M parameters is trained under teacher knowledge distillation along with a small parallel data corpus.
@@
• In-Context Generation. Wang et al. [277] introduced Prompt Diffusion, a novel approach that is jointly trained over multiple tasks using in-context prompts. This method has shown impressive results in high-quality in-context generation for trained tasks and effectively generalizes to new, unseen vision tasks with relevant prompts. Building upon this, Chen et al. [38] further enhance Prompt Diffusion by incorporating a vision encoder-modulated text encoder. This innovation addresses several challenges, including costly pre-training, restrictive problem formulations, limited visual comprehension, and insufficient generalizability to out-of-distribution tasks. Moreover, Najdenkoska et al. [181] propose a novel framework that separates the encoding of the visual context and preserves the structure of the query images. This results in the ability to learn from the visual context and text prompts, but also from either one of them.
@@
• Brain-guided Generation. The brain-guided generation tasks focus on controlling image creation directly from brain activities, such as electroencephalogram (EEG) recordings and functional magnetic resonance imaging (fMRI), bypassing the need to translate thoughts into text. More recently, advancements have been made with the adoption of visual diffusion models, offering enhanced capabilities in accurately translating complex brain activities into coherent visual representations[44, 250, 186, 167, 183, 12, 67].
@@
• Sound-Guided Generation. For sound-guided generation, some works develop an additional audio encoder, utilized to embed input audio into text embedding space or latent feature space to control generation [205, 320, 305]. GlueGen [205] aligns multi-modal encoders such as AudioCLIP with the Stable Diffusion model, enabling sound-to-image generation. Yang et al. [305] propose a unified framework “Align, Adapt, and Inject” (AAI) for sound-guided image generation, editing, and stylization. In particular, this method adapts input sound into a sound token, like an ordinary word, which can plug and play with existing powerful diffusion-based Text-to-Image models.
@@
• Text Rendering. Drawing inspiration from the analysis in unCLIP [213], which highlights the inadequacy of raw CLIP text embeddings in accurately modeling the spelling information in prompts, subsequent efforts such as eDiff-I [13] and Imagen [225] have sought to harness the capabilities of large language models like T5 [209], trained on text-only corpora, as text encoders in image generation. Additionally, DeepFloyd IF, following the design principles of Imagen [225], has demonstrated impressive proficiency in rendering legible text on images, showcasing a significant advancement in this challenging domain. Meanwhile, some approaches are designed to improve text rendering capability for existing text-to-image diffusion models [156, 32, 255, 33, 339, 322, 34, 162]. GlyphControl [306] leverages additional glyph conditional information to enhance the performance of the off-the-shelf Stable-Diffusion model in generating accurate visual text. TextDiffuser et al. [32] first generates the layout of keywords extracted from text prompts and then generates images conditioned on the text prompt and the generated layout. The authors also contribute a large-scale text images dataset with OCR annotations, MARIO-10M, containing 10 million image-text pairs with text recognition, detection, and character-level segmentation annotations. Zhang et al. [322] proposed Diff-Text, a training-free scene text generation framework for any language. Diff-text leverages rendered sketch images as priors to render text by ControlNet [323] and proposes a localized attention constraint to address the unreasonable position problem of scene text.
Figure 5: Illustration of training-free conditional score prediction.
3.1.3 Training-free Conditional Score Prediction
While the above techniques require training, a complementary line of work operates in a training-free fashion (see Fig. 5). These methods analyze and exploit the model’s intrinsic properties so that a pretrained model can directly support novel conditions.
• Attention-map Adjustment. Since attention maps are inherently interpretable, many controllable generation methods build upon attention map adjustment. Attention maps explicitly capture the relationships between tokens, particularly establishing correspondences between image features and textual tokens. This property enables direct control over the placement of textual concepts in generated images, including object positioning (i.e., layout-to-image) [334, 284, 13, 124, 334, 176, 337, 197, 41, 134, 271], fine-grained control of objects and their attributes (i.e., attribute binding) [216], and improving textual concept representation[53].
For instance, since attention mechanisms explicitly model the relationships between text and image tokens, modulating the attention map becomes a pivotal training-free technique for controlling structure in score prediction [13, 124, 334, 176]. eDiff-I [13] presents a technique named “paint-with-words” (also known as pww), rectifying the cross-attention maps of each word by the correspondence segmentation maps to control the location of objects. Additionally, DenseDiffusion [124] introduces a more extensive modulation method by devising multiple regularization, enhancing the precision and flexibility of layout control in score prediction. Furthermore, Chen et al. [41] introduce a soft refinement phase to dismiss the visual boundaries and enhance adjacent interactions.
For the attribute binding task, Ge et al. [71] study the image generation from enriched textual description and propose a region-based diffusion, which constrains the object-level description into the object area via an attention map. Similarly, Structure Diffusion [64] employs linguistic insights to manipulate the cross-attention map, aiming for more accurate attribute binding and improved image composition.
• Feature Injection in Attention. By injecting additional key–value pairs into the attention module, the denoising process can dynamically incorporate visual information from the provided references. [62, 139, 90, 90, 177, 58, 76, 87, 89] Li et al. [139] inject object feature from one of the reference images into the inversion process of another image to realize object placement in self-attention. StyleAligned [89] is designed to produce a series of images that adhere to a given reference style. This method introduces a novel attention sharing mechanism within the self-attention layers, which facilitates the interaction between the features of individual images and those of an additional reference image. Such a design enables the generation process to consider and incorporate style elements from multiple images simultaneously. FreeControl [177] performs PCA to self-attention feature and replaces the principal components of feature in generation by reference components to control object appearance or spatial arrangement. Furthermore, feature injection can also be applied to cross-attention [170, 158]. Pick-and-draw [170] extracts cross-attention map in reference image inversion and generation process, and then injection the inversion feature into the generation process via Earth Movers Distance(EMD) algorithm.
• Others. FreeU et al. [234] enhances diffusion image generation by strategically reweighting skip-connection and backbone feature contributions during inference—boosting coherence and fidelity without any additional training or parameters. Basu et al. [16] propose Mechanistic Localization in text-to-image models, demonstrating that knowledge of visual attributes (e.g., “style,” “objects,” “facts”) can be localized to a small subset of UNet layers, thereby enabling more efficient model editing.
To synthesize high-resolution images, MultiDiffusion [15] formulates an optimization problem that enforces each crop to remain consistent with its denoised counterpart. Although individual denoising steps may introduce conflicting directions, the method fuses them into a unified global denoising step, ultimately producing seamless and high-quality images. Zhou et al. [343] find that the padding is the pivotal mechanism to object arrangement, which is degraded in high-resolution generation. Based on that, they introduce a Progressive Boundary Complement, which creates dynamic virtual image boundaries inside the feature map to enhance position information propagation.
Meanwhile, from the noise initialization perspective, InitNo [77] first samples a lot of initial noise and then designs a cross-attention response score and the self-attention conflict score to evaluate them to find a better one.
Figure 6: Illustration of condition-guided conditional score estimation.
3.2 Condition-guided Score Estimation
While numerous methods adhere to the paradigm of conditional score prediction, some studies explore controllable generation by performing condition prediction from latents or intermediate features during the generative process, and then computing losses against the given conditions to provide gradient guidance for denoising [153, 261, 14, 153, 310, 292, 50, 169], as shown in Fig. 6.
LGP [261] stands as an early pioneer, which innovatively introduces a Latent Edge Predictor, designed to extrapolate sketch information from a series of intermediate features within a UNet architecture. It employs the degree of similarity between the condition sketch and predicted sketch to compute gradients, which are then utilized to guide the score estimation process. Its methodologies and insights have been a source of inspiration for numerous subsequent research endeavors in this field[153, 196, 50, 291]. Furthermore, Universal Guidance [14] and FreeDom [310] are proposed to leverage image-space off-the-shelf predictors to guide denoising. At each denoising step, it attains the clean image by one-step denoising to calculate the guidance gradient.
While the aforementioned methods require a condition predictor to backpropagate condition guidance, layout and segmentation guidance can also be directly estimated through attentiomap, eliminating the need for additional trained models [338, 292, 50, 197, 170, 295, 42, 265, 192]. For instance, BoxDiff [295] designs three spatial constraints (i.e., Inner-Box, Outer-Box, and Corner Constraints) to guide the denoising process. ZestGuide [50] leverages segmentation maps extracted from cross-attention layers, aligning generation with input masks through gradient-based guidance during denoising. To place the object at a specific position, VersaGen [42] calculates the loss from the object-token attention map and the given segmentation. Additionally, VODiff [147] studies the objects’ visibility order and proposes visibility-order-aware loss.
Measuring the representation of textual concepts as a denoising guidance can help improve textual alignment [27, 216, 143]. Attend-and-Excite [27](A&E) represents an early effort in this area, introducing an attention-based Generative Semantic Nursing (GSN) mechanism. This mechanism refines cross-attention units to more effectively ensure that all subjects described in the text prompt are accurately generated. EBAMA [327] extend A&E by introducing an attribute binding loss to address semantic misalignment. Additionally, SynGen [216] employs a unique methodology in text-to-image generation by first conducting a syntactic analysis of the text prompt. This analysis aims to identify entities and their modifiers within the prompt. Following this, SynGen utilizes a novel loss function designed to align the cross-attention maps with the linguistic bindings as indicated by the syntax.
@@ 4 Controllable Generation with Multiple Conditions
The multi-condition generation task aims to generate images under multiple conditions (see Fig. 7). In this section, we conduct a comprehensive overview of these methods from a technical perspective, categorizing them into joint training (Sec. 4.1), continual learning (Sec. 4.2), weight fusion (Sec. 4.3), attention-based integration (Sec. 4.4), and guidance composition (Sec. 4.5). Note that some of the other controllable generation methods also demonstrate multi-condition synthesis capability without dedicated designs [323, 68, 221, 206].
@@ 4.1 Joint Training
Designing a multi-condition framework and jointly training them is a simple yet effective route to realize multi-condition generation. These methods generally focus on multi-condition encoders and training strategies [66, 101, 99, 266, 83, 83, 290, 150, 180, 166].
Composer [101] projects all conditions (including text caption, depthmap, sketch, and etc.) into uniform-dimensional embeddings with the same spatial size as the noisy latent using stacked convolutional layers. It leverages a joint training strategy to generate images from a set of representations, where it uses an independent dropout probability of 0.5 for each condition, a probability of 0.1 for dropping all conditions, and a probability of 0.1 for retaining all conditions. Additionally, Cocktail [99] proposes the controllable normalization method (ControlNorm), which has an additional layer to generate two sets of learnable parameters conditioned on all modalities. These two sets of parameters are used to fuse the external conditional signals and the original signals. UniCombine [266] designs a Conditional MMDiT Attention mechanism, where condition-specific LoRA modules mask the attention across different conditions to naturally support multiple conditions.
@@ 4.2 Continual Learning
Continual learning methods are generally proposed to address knowledge “catastrophic forgetting” in tuning-based conditional score prediction works [154, 236, 246, 235, 80]. Specifically, C-LoRA [236] is composed of a continually self-regularized LoRA in cross-attention layers. It utilizes the past LoRA weight deltas to regulate the new LoRA weight deltas by guiding which parameters are most available to be updated for continual concept learning. Moreover, L2DM [246] devises a task-aware memory enhancement module and an elastic-concept distillation module, which could respectively safeguard the knowledge of both prior concepts and each past personalized concept. It utilizes a rainbow-memory bank strategy to manage long-term and short-term memory and provide regularization samples to safeguard the knowledge in the personalization process. During training, the authors further propose a concept attention artist module and orthogonal attention artist module to update noisec latent for better performance. STAMINA [235] introduces forgetting-regularization and sparsity-regularization in continual learning, avoiding forgetting learned concepts and ensuring no cost to storage or inference. ConceptGuard [80] combines shift embedding, concept-binding prompts, and memory preservation regularization to support new concepts.